1. A 2kg phonograph turntable with diameter 1.5m initially rotating at 4.5 rads makes 4 complete rotations before coming to a stop.

a) What is its angular acceleration
1
Expert's answer
2021-10-10T15:57:04-0400

The angular acceleration is


"\alpha=\frac{\omega^2}{2\phi}=\frac{4.5^2}{2·(4·2\pi)}=0.4\text{ rad/s}^{-1}."